OpenStudy (jamesj):

This question was posted in Mathematics earlier today. An unusual question: in the matrix equation [a b][a]=[20] [c d][e] [43] each variable is positive, an integer, and less than 10. What is a+b+c+d+e?

OpenStudy (dumbcow):

a^2 + be = 20 ac +de = 43 a must be less than 5, If a= 4, Then be = 4 and e must be 1,2,or4 4c + 4d = 43, not possible 43 not divisible by 4 4c + 2d = 43, not possible 43 not divisible by 2 Thus, e = 1 and b = 4 4c + d = 43 ..... c= 9, d = 7 a+b+c+d+e = 4+4+9+7+1 = 25

OpenStudy (asnaseer):

yup - that is pretty much how I solved it as well. 1st equation implies 'a' < 5. 2nd equation implies 'a' and 'e' cannot BOTH be even. so: a = 1 => be = 19 (reject as it implies b or e > 9) a = 2 => be = 16 (reject as 8*2 and 4*4 both lead to 'a' and 'e' being even) a = 3 => be = 11 (reject as it implies b or e > 9) a = 4 => be = 4 (2*2 rejected as 'a' and 'e' would be even, 4*1 accepted) so we get a=4, b=4, e=1 then use 2nd equation with these values to get: 4c + d = 43 max 'd' = 9 => min 'c' = (43-9)/4 = 8.5 min 'd' = 1 => max 'c' = (43-1)/4 = 10.5 therefore c=9 which leads to d=7 so solution is: a=4, b=4, c=9, d=7, e=1 a+b+c+d+e=25

OpenStudy (jamesj):

Yes, asnaseer nailed it when it was originally posted as well. Unusual problem.

OpenStudy (anonymous):

Nice job Asnaseer!

OpenStudy (turingtest):

As James pointed out it's also a great example of the fallibility of Wolfram http://www.wolframalpha.com/input/?i=solve%7B%7B%7Ba%2Cb%7D%2C%7Bc%2Cd%7D%7D.%7B%7Ba%7D%2C%7Be%7D%7D%3D%7B%7B20%7D%2C%7B43%7D%7D%2Ca%2Cb%2Cc%2Cd%2Ce%7D+ Oh how the mighty have fallen :)

OpenStudy (asnaseer):

but do keep in mind that even Wolfram was written by mere mortals :-) we are ALL fallible - you must admit though Wolfram is a very impressive piece of software?

OpenStudy (jamesj):

It is. But I think it's a double-edged sword in the hands of a mathematical beginner.

OpenStudy (asnaseer):

true - especially if you treat it like the "gospel truth". there is no substitute for the human mind (at least not in my opinion)